# Structural Asymmetry and Rigidity of Bounded Solutions for a Parametrized Complex Linear Differential Equation

### v3

@@ 2 Main Proposition
$\delta_{s,s^{\prime}}(u):=x_{s}(u)x_{s^{\prime}}(u)-(x_{s}(u))^{2}=x_{s}(u)(x_{s^{\prime}}(u)-x_{s}(u))>0,\ \forall u\in(t_{*}-\tilde{\epsilon},t_{*}).$
and $\delta_{s,s^{\prime}}(t_{*})=0$. Then we must find $\dot{\delta}_{s,s^{\prime}}(t_{*})\leq 0$. As the previous technique, prove that in our system that $\dot{\delta}_{s,s^{\prime}}(t_{*})>0$. From the differential equation (4), we have
[…]
$x_{s^{\prime}}(t)\dot{x}_{s}(t)$ $=\sigma\,t^{-1}x_{s^{\prime}}(t)x_{s}(t)+\sigma t^{-1}x_{s^{\prime}}(t)+t^{-1}x_{s^{\prime}}(t)\Re\big((1-s)\,t^{-i\tau}\,\eta(t)\big),$
$x_{s}(t)\dot{x}_{s^{\prime}}(t)$ $=\sigma^{\prime}\,t^{-1}x_{s}(t)x_{s^{\prime}}(t)+\sigma^{\prime}t^{-1}x_{s}(t)+t^{-1}x_{s}(t)\Re\big((1-s^{\prime})\,t^{-i\tau}\,\eta(t)\big).$
Implies
$\dot{\delta}_{s,s^{\prime}}(t)$ $=2\sigma\,t^{-1}{\delta}_{s,s^{\prime}}(t)+t^{-1}\big(\sigma x_{s^{\prime}}(t)+\sigma^{\prime}x_{s}(t)\big)$
$+t^{-1}x_{s^{\prime}}(t)\Re\big((1-s)\,t^{-i\tau}\,\eta(t)\big)$
$-t^{-1}x_{s}(t)\Re\big((1-s^{\prime})\,t^{-i\tau}\,\eta(t)\big)$
$-2\sigma t^{-1}x_{s}(t).$
At the point $t_{*}$ we get
$\dot{\delta}_{s,s^{\prime}}(t_{*})$ $=-t_{*}^{-1}\big(\sigma-\sigma^{\prime}\big)x_{s}(t_{*})\big(1-\Re\big(t_{*}^{-i\tau}\,\eta(t_{*})\big)\big).$
Use the fact that $\sigma>\sigma^{\prime}$ and $x_{s}(t_{*})<0$, we obtain $\dot{\delta}_{s,s^{\prime}}(t_{*})>0$. Same, we obtained a contradiction. Then $t_{*}=+\infty$. ∎
3 Proof of the Theorem 4
Intuitively, the solution $\psi_{w}(1,\cdot)$cannot be bounded when $\Re(w)\in\left(\tfrac{1}{2},1\right)$. Indeed, if $\Re(w)\in(0,1)$, then |w|>|1w||w|>|1-w|. Since the initial condition satisfies $\psi_{w}(1,1)=1$ and $\sup_{t\geq 1}|\eta(t)|\leq 1$, for $t$ close to 11 the homogeneous part dominates the non-homogeneous part in differential equation (1). This may therefore lead to the unboundedness of the solution $\psi_{w}(1,\cdot)$. In this manuscript, we compare and study the boundedness of two solutions corresponding to two complex parameters with the same imaginary part, in order to eliminate the imaginary part (which may be very large) from the computations.
Proof of the Theorem 4.
We prove the theorem by contradiction. Suppose there exist $s,s^{\prime}\in B$ with $\Re(s)>\Re(s^{\prime})$ and $\Im(s)=\Im(s^{\prime})$ such that $\big(\mu(s),\,\mu(s^{\prime})\big)=(1,1)$. The solutions $\psi_{s}(1,\cdot)$ and $\psi_{s^{\prime}}(1,\cdot)$ of the differential equation (1) are both bounded on $1,+\infty)$. Thanks to the second Item of the Lemma 3, we have
$\sup_{t\geq 1}\Big|t\Big(\psi_{s}(1,t)+\rho\frac{1-s}{s}\Big)\Big|<+\infty\ \text{and}\ \sup_{t\geq 1}\Big|t\Big(\psi_{s^{\prime}}(1,t)+\rho\frac{1-s^{\prime}}{s^{\prime}}\Big)\Big|<+\infty.$
Use the Proposition 5, we get
$\lim_{t\to+\infty}\Re\Big(t^{-i\Im(s)}\rho\big(\frac{1-s^{\prime}}{s^{\prime}}-\frac{1-s}{s}\big)\Big)\geq 0.$
Since $s,s^{\prime}\in B$, then $\Im(s)\neq 0$, that gives
$\rho\big(\frac{1-s^{\prime}}{s^{\prime}}-\frac{1-s}{s}\big)=0.$
By the hypothesis (H), we have $\rho\neq 0$. then
$\frac{1-s^{\prime}}{s^{\prime}}-\frac{1-s}{s}=0$
We have obtained a contradiction with $\Re(s)\neq\Re(s^{\prime})$. ∎
4 Application to the $\zeta$-Function
The function $\eta(t)=\{t\}$ (the fractional part of $t$) is a locally integrable function and satisfies $0\leq\eta(t)<1$ with a rotation number $\rho$ (here $\rho=1/2$). The Hypothesis (H) is satisfied. By definition of the function $\mu$ in the Definition 2, we have
@@ 4 Application to the $\zeta$-Function
Thanks to the Riemann functional equation [1, Theorem 2.1, p. 13], when $\zeta(s)=0$ then $\zeta(1-\bar{s})=0$. Implies $\big(\mu(s),\ \mu(1-\bar{s})\big)=(1,1)$. By the Theorem 4 we deduce that $\Re(s)=\Re(1-\bar{s})$. The non-trivial zeros of the Riemann zeta function must lie on the critical line.

### v4

@@ 3 Application to the $\zeta$-Function
$\psi_{\frac{1}{2}+i\,\beta}(1,t)-\psi_{i\,\beta}(1,t)$ $\geq t^{\frac{1}{2}}-1-\int_{\lfloor t\rfloor-1}^{t}(t^{\frac{1}{2}}u^{\frac{1}{2}}-u)u^{-2}du$
$+\sum_{n=1}^{\lfloor t\rfloor}\int_{n}^{n+1}\big(n+1+u)\,u^{-2}du$
$-t^{\frac{1}{2}}\sum_{n=1}^{\lfloor t\rfloor}\int_{n}^{n+1}\big((n+1)^{\frac{1}{2}}+u^{\frac{1}{2}}\big)u^{-2}du$
$\geq\frac{1}{3}\big(t^{\frac{1}{2}}-1\big)+\left(1-\frac{1}{t}\right),\qquad\forall t\geq 1.$
We recall that $\sum_{n=1}^{\lfloor t\rfloor-1}J(n+1,n)=0$ for all $t\in[1,2)$. We deduce that
$\Re\Big(\psi_{\frac{1}{2}+i\,\beta}(1,t)-\psi_{i\,\beta}(1,t)\Big)>0,\quad\forall t>1.$
Then the Hypothesis hypothesis H is satisfied. By definition of the function $\mu$ in the Definition 2, we have
@@
Since the Hypothesis hypothesis H is satisfied, by the Theorem 5 when $s\in B$ with $\Re(s)\neq\frac{1}{2}$ we get $\big(\mu(s),\ \mu(1-\bar{s})\big)\neq(1,1)$. The non-trivial zeros of the Riemann zeta function must lie on the critical line.